OpenStudy (anonymous):

Please help me. If cot(theta)=4/3, find csc(theta)

OpenStudy (raden):

does theta in the 1st quadrant ?

OpenStudy (anonymous):

@ZeHanz

OpenStudy (zehanz):

cot(theta)=cos(theta)/sin(theta)=4/3 so cos(theta) = 4*a and sin(theta) = 3*a (see image) If you draw a unit circle, I can only come up with 0.8 and 0.6 for cos and sin (or both negative). This means you can also calculate what csc(theta) is...

OpenStudy (zehanz):

After all, csc(theta) = 1/sin(theta)